OpenStudy (anonymous):

1. What is the first step in solving equations with variables on both sides? A. Add or subtract one of the variable terms from both sides B. Use the distributive property to remove parentheses C. Simplify the equation D. Use the Addition, Subtraction, Multiplication, or Division Properties to solve for the remaining variable 2. Solve: 4x - 7 = 9x + 13 A. x = 0.46 B. x = -7 C. x = 4 D. x = -4 3. Solve: 3x - 5x + 11 = -10 + 2x - 3 A. x = 6 B. x = -6 C. x = -1.1 D. x = 0.5

OpenStudy (anonymous):

3 questions

OpenStudy (anonymous):

lets try to solve prob 2) 4x - 7 = 9x + 13 subtract 4x on both side gives -7 =9x -4x +13 subtract 13 on both sides gives -7-13 =9x-4x now add or subtract like terms gives -20 = 5x divide 5 on both sides gives -20/5=x x=-4

OpenStudy (anonymous):

now this time you can answer prob 1) just understand what process was done on prob 2), also the same as prob 3) the same process you can do... :D